### v1

# Bigger is not Always Better: Scaling Properties of Latent Diffusion Models

Abstract
We study the scaling properties of latent diffusion models (LDMs) with an emphasis on their sampling efficiency. While improved network architecture and inference algorithms have shown to effectively boost sampling efficiency of diffusion models, the role of model size—a critical determinant of sampling efficiency—has not been thoroughly examined. Through empirical analysis of established text-to-image diffusion models, we conduct an in-depth investigation into how model size influences sampling efficiency across varying sampling steps. Our findings unveil a surprising trend: when operating under a given inference budget, smaller models frequently outperform their larger equivalents in generating high-quality results. Moreover, we extend our study to demonstrate the generalizability of the these findings by applying various diffusion samplers, exploring diverse downstream tasks, evaluating post-distilled models, as well as comparing performance relative to training compute. These findings open up new pathways for the development of LDM scaling strategies which can be employed to enhance generative capabilities within limited inference budgets.
Keywords:
Scaling Properties Latent Diffusion Models
1 Introduction
Latent diffusion models (LDMs) [55], and diffusion models in general, trained on large-scale, high-quality data [30, 62] have emerged as a powerful and robust framework for generating impressive results in a variety of tasks, including image synthesis and editing [55, 48, 9, 53, 49], video creation [42, 43, 70, 63], audio production [31], and 3D synthesis [28, 32]. Despite their versatility, the major barrier against wide deployment in real-world applications [12, 8] comes from their low sampling efficiency. The essence of this challenge lies in the inherent reliance of LDMs on multi-step sampling [66, 16] to produce high-quality outputs, where the total cost of sampling is the product of sampling steps and the cost of each step. Specifically, the go-to approach involves using the 50-step DDIM sampling [64, 55], a process that, despite ensuring output quality, still requires a relatively long latency for completion on modern mobile devices with post-quantization. In contrast to single shot generative models (e.g., generative-adversarial networks (GANs) [13]) which bypass the need for iterative refinement [13, 23], the operational latency of LDMs calls for a pressing need for efficiency optimization to further facilitate their practical applications.
Recent advancements in this field [27, 74, 47, 24, 25, 8] have primarily focused on developing faster network architectures with comparable model size to reduce the inference time per step, along with innovations in improving sampling algorithms that allow for using less sampling steps [64, 11, 22, 34, 33, 71]. Further progress has been made through diffusion-distillation techniques [36, 59, 65, 61, 14, 41], which simplifies the process by learning multi-step sampling results in a single forward pass, and then broadcasts this single-step prediction multiple times. These distillation techniques leverage the redundant learning capability in LDMs, enabling the distilled models to assimilate additional distillation knowledge. Despite these efforts being made to improve diffusion models, the sampling efficiency of smaller, less redundant models has not received adequate attention. A significant barrier to this area of research is the scarcity of available modern accelerator clusters [20], as training high-quality text-to-image (T2I) LDMs from scratch is both time-consuming and expensive—often requiring several weeks and hundreds of thousands of dollars.
In this paper, we empirically investigate the scaling properties of LDMs, with a particular focus on understanding how their scaling properties impact the sampling efficiency across various model sizes. We trained a suite of 12 text-to-image LDMs from scratch, ranging from 39 million to 5 billion parameters, under a constrained budget. Example results are depicted in Fig. 1. All models were trained on TPUv5 using internal data sources with about 600 million aesthetically-filtered text-to-image pairs. Our study reveals that there exist a scaling trend within LDMs, notably that smaller models may have the capability to surpass larger models under an equivalent sampling budget. Furthermore, we investigate how the size of pre-trained text-to-image LDMs affects their sampling efficiency across diverse downstream tasks, such as real-world super-resolution [58, 57] and subject-driven text-to-image synthesis (i.e., Dreambooth) [56].
1.1 Summary
Our key findings for scaling latent diffusion models in text-to-image generation and various downstream tasks are as follows:
Pretraining performance scales with training compute. We demonstrate a clear link between compute resources and LDM performance by scaling models from 39 million to 5 billion parameters. This suggests potential for further improvement with increased scaling. See Section 3.1 for details.
Downstream performance scales with pretraining. We demonstrate a strong correlation between pretraining performance and success in downstream tasks. Smaller models, even with extra training, cannot fully bridge the gap created by the pretraining quality of larger models. This is explored in detail in Section 3.2.
Smaller models sample more efficient. Smaller models initially outperform larger models in image quality for a given sampling budget, but larger models surpass them in detail generation when computational constraints are relaxed. This is further elaborated in Section 3.3.1 and Section 3.3.2.
Sampler does not change the scaling efficiency. Smaller models consistently demonstrate superior sampling efficiency, regardless of the diffusion sampler used. This holds true for deterministic DDIM [64], stochastic DDPM [16], and higher-order DPM-Solver++ [35]. For more details, see Section 3.4.
Smaller models sample more efficient on the downstream tasks with fewer steps. The advantage of smaller models in terms of sampling efficiency extends to the downstream tasks when using less than 20 sampling steps. This is further elaborated in Section 3.5.
Diffusion distillation does not change scaling trends. Even with diffusion distillation, smaller models maintain competitive performance against larger distilled models when sampling budgets are constrained. This suggests distillation does not fundamentally alter scaling trends. See Section 3.6 for in-depth analysis.
2 Related Work
Scaling laws.
Recent Large Language Models (LLMs) including GPT [3], PaLM [2], and LLaMa [67] have dominated language generative modeling tasks. The foundational works [21, 3, 18] for investigating their scaling behavior have shown the capability of predicting the performance from the model size. They also investigated the factors that affect the scaling properties of language models, including training compute, dataset size and quality, learning rate schedule, etc. Those experimental clues have effectively guided the later language model development, which have led to the emergence of several parameter-efficient LLMs [18, 67, 75, 1]. However, scaling generative text-to-image models are relatively unexplored, and existing efforts have only investigated the scaling properties on small datasets or small models, like scaling UNet [46] to 270 million parameters and DiT [47] on ImageNet (14 million), or less-efficient autoregressive models [6]. Different from these attempts, our work investigates the scaling properties by scaling down the efficient and capable diffusion models, i.e. LDMs [55], on internal data sources that have about 600 million aesthetics-filtered text-to-image pairs for featuring the sampling efficiency of scaled LDMs. We also scale LDMs on various scenarios such as finetuning LDMs on downstream tasks [69, 56] and distilling LDMs [41] for faster sampling to demonstrate the generalizability of the scaled sampling-efficiency.
Efficient diffusion models.
Nichol et al. [46] show that the generative performance of diffusion models improves as the model size increases. Based on this preliminary observation, the model size of widely used LDMs, e.g., Stable Diffusion [55], has been empirically increased to billions of parameters [51, 48]. However, such a large model makes it impossible to fit into the common inference budget of practical scenarios. Recent work on improving the sampling efficiency focus on improving network architectures [27, 74, 47, 24, 25, 8] or the sampling procedures [64, 11, 22, 34, 33, 71]. We explore sampling efficiency by training smaller, more compact LDMs. Our analysis involves scaling down the model size, training from scratch, and comparing performance at equivalent inference cost.
(a) 39M model
(b) 83M model
(c) 145M model
(d) 223M model
(e) 318M model
(f) 430M model
(g) 558M model
(h) 704M model
(i) 2B model
Figure 1: Text-to-image results from our scaled LDMs (39M - 2B), highlighting the improvement in visual quality with increased model size (note: 39M model is the exception). All images generated using 50-step DDIM sampling. We use representative prompts from PartiPrompts [72], including “a professional photo of a sunset behind the grand canyon.”, “Dogs sitting around a poker table with beer bottles and chips. Their hands are holding cards.”, ‘Portrait of anime girl in mechanic armor in night Tokyo.”, “a teddy bear on a skateboard.”, “a pixel art corgi pizza.”, “Snow mountain and tree reflection in the lake.”, “a propaganda poster depicting a cat dressed as french emperor napoleon holding a piece of cheese.”, “a store front that has the word ‘LDMs’ written on it.”, and “ten red apples.”. Check our supplement for additional visual comparisons.
Params 39M 83M 145M 223M 318M 430M 558M 704M 866M 2B 5B
Filters $(c)$ 64 96 128 160 192 224 256 288 320 512 768
GFLOPS 25.3 102.7 161.5 233.5 318.5 416.6 527.8 652.0 789.3 1887.5 4082.6
Norm. Cost 0.07 0.13 0.20 0.30 0.40 0.53 0.67 0.83 1.00 2.39 5.17
FID $\downarrow$ 25.30 24.30 24.18 23.76 22.83 22.35 22.15 21.82 21.55 20.98 20.14
CLIP $\uparrow$ 0.305 0.308 0.310 0.310 0.311 0.312 0.312 0.312 0.312 0.312 0.314
Table 1: We scale the baseline LDM (i.e., 866M Stable Diffusion v1.5) by changing the base number of channels $c$ that controls the rest of the U-Net architecture as $[c,2c,4c,4c]$ (See Fig. 2). GFLOPS are measured for an input latent of shape $64\times 64\times 4$ with FP32. We also show a normalized running cost with respect to the baseline model. The text-to-image performance (FID and CLIP scores) for our scaled LDMs is evaluated on the COCO-2014 validation set with 30k samples, using 50-step DDIM sampling. It is worth noting that all the model sizes, and the training and the inference costs reported in this work only refer to the denoising UNet in the latent space, and do not include the 1.4B text encoder and the 250M latent encoder and decoder.
Efficient non-diffusion generative models.
Compared to diffusion models, other generative models such as, Variational Autoencoders (VAEs) [26, 54, 39, 68], Generative Adversarial Networks (GANs) [13, 40, 23, 52, 45], and Masked Models [10, 50, 15, 5, 4], are more efficient, as they rely less on an iterative refinement process. Sauer et al. [60] recently scaled up StyleGAN [23] into 1 billion parameters and demonstrated the single-step GANs’ effectiveness in modeling text-to-image generation. Chang et al. [4] scaled up masked transformer models for text-to-image generation. These non-diffusion generative models can generate high-quality images with less inference cost, which require fewer sampling steps than diffusion models and autoregressive models, but they need more parameters, i.e., 4 billion parameters.
3 Scaling LDMs
We developed a family of powerful Latent Diffusion Models (LDMs) built upon the widely-used 866M Stable Diffusion v1.5 standard [55] [^1] We adopted SD v1.5 since it is among the most popular diffusion models https://huggingface.co/models?sort=likes.. The denoising UNet of our models offers a flexible range of sizes, with parameters spanning from 39M to 5B. We incrementally increase the number of filters in the residual blocks while maintaining other architecture elements the same, enabling a predictably controlled scaling. Table 1 shows the architectural differences among our scaled models. We also provide the relative cost of each model against the baseline model. Fig. 2 shows the architectural differences during scaling. Models were trained using internal data sources with about 600 million aesthetically-filtered text-to-image pairs. All the models are trained for 500K steps, batch size 2048, and learning rate 1e-4. This allows for all the models to have reached a point where we observe diminishing returns. Fig. 1 demonstrates the consistent generation capabilities across our scaled models. We used the common practice of 50 sampling steps with the DDIM sampler, 7.5 classifier-free guidance rate, for text-to-image generation. The visual quality of the results exhibits a clear improvement as model size increases.
In order to evaluate the performance of the scaled models, we test the text-to-image performance of scaled models on the validation set of COCO 2014 [30] with 30k samples. For downstream performance, specifically real-world super-resolution, we test the performance of scaled models on the validation of DIV2K with 3k randomly cropped patches, which are degraded with the RealESRGAN degradation [69].
Figure 2: Our scaled latent diffusion models vary in the number of filters within the denoising U-Net. Other modules remain consistent. Smooth channel scaling (64 to 768) within residual blocks yields models ranging from 39M to 5B parameters. For downstream tasks requiring image input, we use an encoder to generate a latent code; this code is then concatenated with the noise vector in the denoising U-Net.
Refer to captionRefer to caption
Figure 3: In text-to-image generation, we observe consistent trends across various model sizes in how quality metrics (FID and CLIP scores) relate to training compute (i.e., the product of training steps and model GFLOPS). Under moderate training resources ($<1G$) training compute is the most relevant factor dominating quality. When training compute is unlimited (e.g., $\geq 1G$), larger models ($>700M$) lead to superior performance. Note that the 5B model’s size exceeds single TPU capacity, forcing memory-efficient training. This regime has different learning dynamics than smaller models (e.g., 2B), impacting accuracy at comparable training compute.
3.1 Training compute scales text-to-image performance
Refer to captionRefer to caption
Figure 4: In 4×4$\times$4 × real image super-resolution, FID and LPIPS scores reveal an interesting divergence. Model size drives FID improvement, while training compute most impacts LPIPS. Despite this, visual assessment (Fig. 5) confirms the importance of model size for superior detail recovery (similarly as observed in the text-to-image pretraining).
We find that our scaled LDMs, across various model sizes, exhibit similar trends in generative performance relative to training compute cost, especially after training stabilizes, which typically occurs after 200K iterations. These trends demonstrate a smooth scaling in learning capability between different model sizes. To elaborate, Fig. 3 illustrates a series of training runs with models varying in size from 39 million to 5 billion parameters, where the training compute cost is quantified as the product of relative cost shown in Table 1 and training iterations. Model performance is evaluated by using the same sampling steps and sampling parameters. In scenarios with moderate training compute (i.e., $<1G$, see Fig. 3), the generative performance of T2I models scales well with additional compute resources.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
83M 145M 223M 318M 430M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LR 558M 704M 866M 2B HR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
83M 145M 223M 318M 430M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LR 558M 704M 866M 2B HR
Figure 5: In 4$\times$ super-resolution, visual quality directly improves with increased model size. As these scaled models vary in pretraining performance, the results clearly demonstrate that pretraining boosts super-resolution capabilities in both quantitative (Fig 4) and qualitative ways. Additional results are given in supplementary material.
Refer to caption Refer to caption Refer to caption Refer to caption
83M 145M 223M 318M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Inputs 430M 558M 866M 2B
Refer to caption Refer to caption Refer to caption Refer to caption
83M 145M 223M 318M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Inputs 430M 558M 866M 2B
Figure 6: Visualization of the Dreambooth results shows two distinct tiers based on model size. Smaller models (83M-223M) perform similarly, as do larger ones (318M-2B), with a clear quality advantage for the larger group. Additional results are given in supplementary material.
3.2 Pretraining scales downstream performance
Using scaled models based on their pretraining on text-to-image data, we finetune these models on the downstream tasks of real-world super-resolution [58, 57] and DreamBooth [56]. The performance of these pretrained models is shown in Table. 1. In the left panel of Fig. 4, we present the generative performance FID versus training compute on the super-resolution (SR) task. It can be seen that the performance of SR models is more dependent on the model size than training compute. Our results demonstrate a clear limitation of smaller models: they cannot reach the same performance levels as larger models, regardless of training compute.
While the distortion metric LPIPS shows some inconsistencies compared to the generative metric FID (Fig. 4), Fig. 5 clearly demonstrates that larger models excel in recovering fine-grained details compared to smaller models.
The key takeaway from Fig. 4 is that large super-resolution models achieve superior results even after short finetuning periods compared to smaller models. This suggests that pretraining performance (dominated by the pretraining model sizes) has a greater influence on the super-resolution FID scores than the duration of finetuning (i.e., training compute for finetuning).
Furthermore, we compare the visual results of the DreamBooth finetuning on the different models in Fig. 6. We observe a similar trend between visual quality and model size. Please see our supplement for more discussions on the other quality metrics.
3.3 Scaling sampling-efficiency
(a) 50-step sampling results of the 145M model
(b) 50-step sampling results of the 866M model
Figure 7: Visualization of text-to-image results with different CFG rates (from left to right in each row: $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$). The prompt used is “A raccoon wearing formal clothes, wearing a top hat and holding a cane. Oil painting in the style of Rembrandt.”. We observe that changes in CFG rates impact visual quality more significantly than the prompt semantic accuracy. We use the FID score for quantitative determination of optimal sampling performance (Fig. 8) because it directly measures visual quality, unlike the CLIP score, which focuses on semantic similarity.
Figure 8: The impact of the CFG rate on text-to-image generation depends on the model size and sampling steps. As demonstrated in the left and center panels, the optimal CFG rate changes as the sampling steps increased. To determine the optimal performance (according to the FID score) of each model and each sampling steps, we systematically sample the model at various CFG rates and identify the best one. As a reference of the optimal performance, the right panel shows the CFG rate corresponding to the optimal performance of each model for a given number of sampling steps.
3.3.1 Analyzing the effect of CFG rate.
Text-to-image generative models require nuanced evaluation beyond single metrics. Sampling parameters are vital for customization, with the Classifier-Free Guidance (CFG) rate [17] directly influencing the balance between visual fidelity and semantic alignment with text prompt. Rombach et al. [55] experimentally demonstrate that different CFG rates result in different CLIP and FID scores.
Refer to captionRefer to caption
Figure 9: Comparison of text-to-image performance of models with varying sizes. The left figure shows the relationship between sampling cost (normalized cost $\times$ sampling steps) and sampling steps for different model sizes. The right figure plots the text-to-image FID score as a function of the sampling cost for the same models. Key Observation: Smaller models achieve better FID scores than larger models for a fixed sampling cost. For instance, at a cost of 3, the 83M model achieves the best FID compared to the larger models. This suggests that smaller models might be more efficient in achieving good visual quality with lower costs. We highlight the best performing model at different sampling costs [3–48].
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(a) Prompt: “A corgi’s head depicted as a nebula.”. Sampling Cost $\approx$ 6.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(b) Prompt: “A pineapple surfing on a wave.”. Sampling Cost $\approx$ 12.
Figure 10: Text-to-image results of the scaled LDMs under approximately the same inference cost (normalized cost $\times$ sampling steps). Smaller models can produce comparable or even better visual results than larger models under similar sampling cost.
Refer to captionRefer to caption
Figure 11: Left: Text-to-image performance FID as a function of the sampling cost (normalized cost $\times$ sampling steps) for the DDPM sampler (solid curves) and the DDIM sampler (dashed curves). Right: Text-to-image performance FID as a function of the sampling cost for the second-order DPM-Solver++ sampler (solid curves) and the DDIM sampler (dashed curves). Suggested by the trends shown in Fig. 9, we only show the sampling steps $\leq 50$ as using more steps does not improve the performance.
In this study, we find that CFG rate as a sampling parameter yields inconsistent results across different model sizes. Hence, it is interesting to quantitatively determine the optimal CFG rate for each model size and sampling steps using either FID or CLIP score. We demonstrate this by sampling the scaled models using different CFG rates, i.e., $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$ and comparing their quantitative and qualitative results. In Fig. 7, we present visual results of two models under varying CFG rates, highlighting the impact on the visual quality. We observed that changes in CFG rates impact visual quality more significantly than prompt semantic accuracy and therefore opted to use the FID score for quantitative determination of the optimal CFG rate. performance. Fig. 8 shows how different classifier-free guidance rates affect the FID scores in text-to-image generation (see figure caption for more details).
3.3.2 Scaling efficiency trends.
Using the optimal CFG rates established for each model at various number of sampling steps, we analyze the optimal performance to understand the sampling efficiency of different LDM sizes. Specifically, in Fig. 9, we present a comparison between different models and their optimal performance given the sampling cost (normalized cost $\times$ sampling steps). By tracing the points of optimal performance across various sampling cost—represented by the dashed vertical line—we observe a consistent trend: smaller models frequently outperform larger models across a range of sampling cost in terms of FID scores. Furthermore, to visually substantiate better-quality results generated by smaller models against larger ones, Fig. 10 compares the results of different scaled models, which highlights that the performance of smaller models can indeed match their larger counterparts under similar sampling cost conditions. Please see our supplement for more visual comparisons.
3.4 Scaling sampling-efficiency in different samplers
To assess the generalizability of observed scaling trends in sampling efficiency, we compared scaled LDM performance using different diffusion samplers. In addition to the default DDIM sampler, we employed two representative alternatives: the stochastic DDPM sampler [16] and the high-order DPM-Solver++ [35].
Experiments illustrated in Fig. 11 reveal that the DDPM sampler typically produces lower-quality results than DDIM with fewer sampling steps, while the DPM-Solver++ sampler generally outperforms DDIM in image quality (see the figure caption for details). Importantly, we observe consistent sampling-efficiency trends with the DDPM and DPM-Solver++ sampler as seen with the default DDIM: smaller models tend to achieve better performance than larger models under the same sampling cost. Since the DPM-Solver++ sampler is not designed for use beyond 20 steps, we focused our testing within this range. This finding demonstrates that the scaling properties of LDMs remain consistent regardless of the diffusion sampler used.
3.5 Scaling downstream sampling-efficiency
Refer to captionRefer to caption
Figure 12: Super-resolution performance vs. sampling cost for different model sizes. Left: FID scores of super-resolution models under limited sampling steps (less than or equal to 20). Smaller models tend to achieve lower (better) FID scores within this range. Right: FID scores of super-resolution models under a larger number of sampling steps (greater than 20). Performance differences between models become less pronounced as sampling steps increase.
Refer to captionRefer to caption
Figure 13: Distillation improves text-to-image performance and scalability. Left: Distilled Latent Diffusion Models (LDMs) consistently exhibit lower (better) FID scores compared to their undistilled counterparts across varying model sizes. The consistent acceleration factor (approx. 5×5$\times$5 ×) indicates that the benefits of distillation scale well with model size. Right: Distilled models using only 4 sampling steps achieve FID scores comparable to undistilled models using significantly more steps. Interestingly, at a sampling cost of 7, the distilled 866M model performs similarly to the smaller, undistilled 83M model, suggesting improved efficiency. See the supplement for visual examples demonstrating the quality improvements achieved through distillation.
Here, we investigate the scaling sampling-efficiency of LDMs on downstream tasks, specifically focusing on the super-resolution task. Unlike our earlier discussions on optimal sampling performance, there is limited literature demonstrating the positive impacts of SR performance without using classifier-free guidance. Thus, our approach directly uses the SR sampling result without applying classifier-free guidance. Inspired from Fig. 4, where the scaled downstream LDMs have significant performance difference in 50-step sampling, we investigate sampling efficiency from two different aspects, i.e., fewer sampling steps $[4,20]$ and more sampling steps $(20,250]$. As shown in the left part of Fig. 12, the scaling sampling-efficiency still holds in the SR tasks when the number of sampling steps is less than or equal to 20 steps. Beyond this threshold, however, larger models demonstrate greater sampling-efficiency than smaller models, as illustrated in the right part of Fig. 12. This observation suggests the consistent sampling efficiency of scaled models on fewer sampling steps from text-to-image generation to super-resolution tasks.
3.6 Scaling sampling-efficiency in distilled LDMs.
We have featured the scaling sampling-efficiency of latent diffusion models, which demonstrates that smaller model sizes exhibit higher sampling efficiency. A notable caveat, however, is that smaller models typically imply reduced modeling capability. This poses a challenge for recent diffusion distillation methods [36, 59, 65, 61, 14, 41, 37, 29] that heavily depend on modeling capability. One might expect a contradictory conclusion and believe the distilled large models sample faster than distilled small models. In order to demonstrate the sampling efficiency of scaled models after distillation, we distill our previously scaled models with conditional consistency distillation [65, 41] on text-to-image data and compare those distilled models on their optimal performance. Please see our supplement for more distillation details.
To elaborate, we test all distilled models with the same 4-step sampling, which is shown to be able to achieve the best sampling performance; we then compare each distilled model with the undistilled one on the normalized sampling cost. We follow the same practice discussed in Section 3.3.1 for selecting the optimal CFG rate and compare them under the same relative inference cost. The results shown in the left part of Fig. 13 demonstrate that distillation significantly improves the generative performance for all models in 4-step sampling, with FID improvements across the board. By comparing these distilled models with the undistilled models in the right part of Fig. 13, we demonstrate that distilled models outperform undistilled models at the same sampling cost. However, at the specific sampling cost, i.e., sampling cost $\approx\texttt{8}$, the smaller undistilled 83M model still achieves similar performance to the larger distilled 866M model. The observation further supports our proposed scaling sampling-efficiency of LDMs, which still holds under the circumstance of diffusion distillation.
4 Conclusion
In this paper, we investigated scaling properties of Latent Diffusion Models (LDMs), specifically through scaling model size from 39 million to 5 billion parameters. We trained these scaled models from scratch on a large-scale text-to-image dataset and then finetuned the pretrained models for downstream tasks. Our findings unveil that, under identical sampling costs, smaller models frequently outperform larger models, suggesting a promising direction for accelerating LDMs in terms of model size. We further show that the sampling efficiency is consistent in multiple axes. For example, it is invariant to various diffusion samplers (stochastic and deterministic), and also holds true for distilled models. We believe this analysis of scaling sampling efficiency would be instrumental in guiding future developments of LDMs, specifically for balancing model size against performance and efficiency in a broad spectrum of practical applications.
Limitations.
This work utilizes visual quality inspection alongside established metrics like FID and CLIP scores. We opted to avoid human evaluations due to the immense number of different combinations needed for the more than 1000 variants considered in this study. However, it is important to acknowledge the potential discrepancy between visual quality and quantitative metrics, which is actively discussed in recent works [73, 19, 7]. Furthermore, claims regarding the scalability of latent models are made specifically for the particular model family studied in this work. Extending this analysis to other model families, particularly those incorporating transformer-based backbones [47, 43, 38], would be a valuable direction for future research.
5 Acknowledgments
We are grateful to Keren Ye, Jason Baldridge for their valuable feedback. We also extend our gratitude to Shlomi Fruchter, Kevin Murphy, Mohammad Babaeizadeh, and Han Zhang for their instrumental contributions in facilitating the initial implementation of the latent diffusion models.
Appendix 0.A Scaling the text-to-image performance
In order to provide detailed visual comparisons for Fig. 1 in the main manuscript, Fig. 14, Fig. 15, and Fig. 16 show the generated results with the same prompt and the same sampling parameters (i.e., 50-step DDIM sampling and 7.5 CFG rate).
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(a) Prompt: “a professional photo of a sunset behind the grand canyon.”
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(b) Prompt: “Dogs sitting around a poker table with beer bottles and chips. Their hands are holding cards.”
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(c) Prompt: ‘Portrait of anime girl in mechanic armor in night Tokyo.”
Figure 14: Text-to-image results from our scaled LDMs (83M - 5B), highlighting the improvement in visual quality with increased model size.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(a) Prompt: “a teddy bear on a skateboard.”
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(b) Prompt: “a pixel art corgi pizza.”
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(c) Prompt: “Snow mountain and tree reflection in the lake.”
Figure 15: Text-to-image results from our scaled LDMs (83M - 5B), highlighting the improvement in visual quality with increased model size.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(a) Prompt: “a propaganda poster depicting a cat dressed as french emperor napoleon holding a piece of cheese.”
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(b) Prompt: “a store front that has the word ‘LDMs’ written on it.”
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
Refer to caption 2B
Refer to caption 5B
(c) Prompt: “ten red apples.”
Figure 16: Text-to-image results from our scaled LDMs (83M - 5B), highlighting the improvement in visual quality with increased model size.
Appendix 0.B Scaling downstream performance
To provide more metrics for the super-resolution experiments in Fig. 4 of the main manuscript, Fig. 18 shows the generative metric IS for the super-resolution results. Fig. 18 shows the visual results of the super-resolution results in order to provide more visual results for the visual comparisons of Fig. 5 in the main manuscript.
Refer to captionRefer to caption
Figure 17: For super-resolution, we show the trends between the generative metric IS and the training compute still depend on the pretraining, which is similar to the trends between the generative metric FID and the training compute.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
83M 145M 223M 318M 430M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LR 558M 704M 866M 2B HR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
83M 145M 223M 318M 430M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LR 558M 704M 866M 2B HR
Figure 18: In 4$\times$ super-resolution, visual quality directly improves with increased model size. As these scaled models vary in pretraining performance, the results clearly demonstrate that pretraining boosts super-resolution capabilities.
Appendix 0.C Scaling sampling-efficiency in distilled LDMs
Diffusion distillation methods for accelerating sampling are generally derived from Progressive Distillation (PD) [59] and Consistency Models (CM) [65]. In the main paper, we have shown that CoDi [41] based on CM is scalable to different model sizes. Here we show other investigated methods, i.e., guided distillation [44], has inconsistent acceleration effects across different model sizes. Fig. 19 shows guided distillation results for the 83M and 223M models respectively, where s16 and s8 denote different distillation stages. It is easy to see that the performance improvement of these two models is inconsistent.
Fig. 20 shows the visual results of the CoDi distilled models and the undistilled models under the same sampling cost to demonstrate the sampling-efficiency.
Refer to captionRefer to caption
Figure 19: Left: Guided distillation on the 83M model for text-to-image generation. Right: Guided distillation on the 224M model for text-to-image generation.
Refer to caption 866M 4-step distilled
Refer to caption 866M 4-step undistilled
Refer to caption 83M 50-step undistilled
(a) Prompt: “A corgi’s head depicted as a nebula.”. Sampling Cost $\approx$ 7.
Refer to caption 866M 4-step distilled
Refer to caption 866M 4-step undistilled
Refer to caption 83M 50-step undistilled
(b) Prompt: “a cat drinking a pint of beer.”. Sampling Cost $\approx$ 7.
Figure 20: We visualize text-to-image generation results of the tested LDMs under approximately the same inference cost.
Appendix 0.D Scaling the sampling-efficiency
To provide more visual comparisons additional to Fig. 10 in the main paper, Fig. 21, Fig.22, and Fig. 23 present visual comparisons between different scaled models under a uniform sampling cost. This highlights that the performance of smaller models can indeed match their larger counterparts under similar sampling cost.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(a) Prompt: “a cat drinking a pint of beer.”. Sampling Cost $\approx$ 10.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(b) Prompt: “a teddy bear on a skateboard.”. Sampling Cost $\approx$ 10.
Figure 21: We visualize text-to-image generation results of the tested LDMs under approximately the same inference cost. We observe that smaller models can produce comparable or even better visual results than larger models under similar sampling cost (model GFLOPs $\times$ sampling steps).
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(a) Prompt: “a propaganda poster depicting a cat dressed as french emperor napoleon holding a piece of cheese.”. Sampling Cost $\approx$ 14.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(b) Prompt: “Dogs sitting around a poker table with beer bottles and chips.”. Sampling Cost $\approx$ 14.
Figure 22: We visualize text-to-image generation results of the tested LDMs under approximately the same inference cost. We observe that smaller models can produce comparable or even better visual results than larger models under similar sampling cost (model GFLOPs $\times$ sampling steps).
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(a) Prompt: “a pixel art corgi pizza.”. Sampling Cost $\approx$ 18.
Refer to caption 83M
Refer to caption 145M
Refer to caption 223M
Refer to caption 318M
Refer to caption 430M
Refer to caption 558M
Refer to caption 704M
Refer to caption 866M
(b) Prompt: “Snow mountain and tree reflection in the lake.”. Sampling Cost $\approx$ 18.
Figure 23: We visualize text-to-image generation results of the tested LDMs under approximately the same inference cost. We observe that smaller models can produce comparable or even better visual results than larger models under similar sampling cost (model GFLOPs $\times$ sampling steps).

### v2

@@
 These findings open up new pathways for the development of LDM scaling strategies which can be employed to enhance generative capabilities within limited inference budgets.
-Keywords:
-Scaling Properties Latent Diffusion Models
 1 Introduction
-Latent diffusion models (LDMs) [55], and diffusion models in general, trained on large-scale, high-quality data [30, 62] have emerged as a powerful and robust framework for generating impressive results in a variety of tasks, including image synthesis and editing [55, 48, 9, 53, 49], video creation [42, 43, 70, 63], audio production [31], and 3D synthesis [28, 32].
-Despite their versatility, the major barrier against wide deployment in real-world applications [12, 8] comes from their low sampling efficiency.
-The essence of this challenge lies in the inherent reliance of LDMs on multi-step sampling [66, 16] to produce high-quality outputs, where the total cost of sampling is the product of sampling steps and the cost of each step.
-Specifically, the go-to approach involves using the 50-step DDIM sampling [64, 55], a process that, despite ensuring output quality, still requires a relatively long latency for completion on modern mobile devices with post-quantization.
-In contrast to single shot generative models (e.g., generative-adversarial networks (GANs) [13]) which bypass the need for iterative refinement [13, 23], the operational latency of LDMs calls for a pressing need for efficiency optimization to further facilitate their practical applications.
-Recent advancements in this field [27, 74, 47, 24, 25, 8] have primarily focused on developing faster network architectures with comparable model size to reduce the inference time per step, along with innovations in improving sampling algorithms that allow for using less sampling steps [64, 11, 22, 34, 33, 71].
-Further progress has been made through diffusion-distillation techniques [36, 59, 65, 61, 14, 41], which simplifies the process by learning multi-step sampling results in a single forward pass, and then broadcasts this single-step prediction multiple times.
+Latent diffusion models (LDMs) (Rombach et al., 2022), and diffusion models in general, trained on large-scale, high-quality data (Lin et al., 2014; Schuhmann et al., 2022) have emerged as a powerful and robust framework for generating impressive results in a variety of tasks, including image synthesis and editing (Rombach et al., 2022; Podell et al., 2023; Delbracio & Milanfar, 2023; Ren et al., 2023; Qi et al., 2023), video creation (Mei & Patel, 2023; Mei et al., 2023; Wu et al., 2023; Singer et al., 2022), audio production (Liu et al., 2023a), and 3D synthesis (Lin et al., 2023; Liu et al., 2023b).
+Despite their versatility, the major barrier against wide deployment in real-world applications (Du et al., 2023; Choi et al., 2023) comes from their low sampling efficiency.
+The essence of this challenge lies in the inherent reliance of LDMs on multi-step sampling (Song et al., 2021b; Ho et al., 2020) to produce high-quality outputs, where the total cost of sampling is the product of sampling steps and the cost of each step.
+Specifically, the go-to approach involves using the 50-step DDIM sampling (Song et al., 2021a; Rombach et al., 2022), a process that, despite ensuring output quality, still requires a relatively long latency for completion on modern mobile devices with post-quantization.
+In contrast to single shot generative models (e.g., generative-adversarial networks (GANs) (Goodfellow et al., 2020)) which bypass the need for iterative refinement (Goodfellow et al., 2020; Karras et al., 2019), the operational latency of LDMs calls for a pressing need for efficiency optimization to further facilitate their practical applications.
+Recent advancements in this field (Li et al., 2023; Zhao et al., 2023; Peebles & Xie, 2023; Kim et al., 2023b; a; Choi et al., 2023) have primarily focused on developing faster network architectures with comparable model size to reduce the inference time per step, along with innovations in improving sampling algorithms that allow for using less sampling steps (Song et al., 2021a; Dockhorn et al., 2022; Karras et al., 2022; Lu et al., 2022a; Liu et al., 2023c; Xu et al., 2023).
+Further progress has been made through diffusion-distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Song et al., 2023; Sauer et al., 2023b; Gu et al., 2023; Mei et al., 2024a), which simplifies the process by learning multi-step sampling results in a single forward pass, and then broadcasts this single-step prediction multiple times.
 These distillation techniques leverage the redundant learning capability in LDMs, enabling the distilled models to assimilate additional distillation knowledge.
 Despite these efforts being made to improve diffusion models, the sampling efficiency of smaller, less redundant models has not received adequate attention.
-A significant barrier to this area of research is the scarcity of available modern accelerator clusters [20], as training high-quality text-to-image (T2I) LDMs from scratch is both time-consuming and expensive—often requiring several weeks and hundreds of thousands of dollars.
+A significant barrier to this area of research is the scarcity of available modern accelerator clusters (Jouppi et al., 2023), as training high-quality text-to-image (T2I) LDMs from scratch is both time-consuming and expensive—often requiring several weeks and hundreds of thousands of dollars.
 In this paper, we empirically investigate the scaling properties of LDMs, with a particular focus on understanding how their scaling properties impact the sampling efficiency across various model sizes.
@@
 Our study reveals that there exist a scaling trend within LDMs, notably that smaller models may have the capability to surpass larger models under an equivalent sampling budget.
-Furthermore, we investigate how the size of pre-trained text-to-image LDMs affects their sampling efficiency across diverse downstream tasks, such as real-world super-resolution [58, 57] and subject-driven text-to-image synthesis (i.e., Dreambooth) [56].
+Furthermore, we investigate how the size of pre-trained text-to-image LDMs affects their sampling efficiency across diverse downstream tasks, such as real-world super-resolution (Saharia et al., 2022; Sahak et al., 2023) and subject-driven text-to-image synthesis (i.e., Dreambooth) (Ruiz et al., 2023).
 1.1 Summary
@@
 Smaller models consistently demonstrate superior sampling efficiency, regardless of the diffusion sampler used.
-This holds true for deterministic DDIM [64], stochastic DDPM [16], and higher-order DPM-Solver++ [35].
+This holds true for deterministic DDIM (Song et al., 2021a), stochastic DDPM (Ho et al., 2020), and higher-order DPM-Solver++ (Lu et al., 2022b).
 For more details, see Section 3.4.
@@
 Scaling laws.
-Recent Large Language Models (LLMs) including GPT [3], PaLM [2], and LLaMa [67] have dominated language generative modeling tasks.
-The foundational works [21, 3, 18] for investigating their scaling behavior have shown the capability of predicting the performance from the model size.
+Recent Large Language Models (LLMs) including GPT (Brown et al., 2020), PaLM (Anil et al., 2023), and LLaMa (Touvron et al., 2023) have dominated language generative modeling tasks.
+The foundational works (Kaplan et al., 2020; Brown et al., 2020; Hoffmann et al., 2022) for investigating their scaling behavior have shown the capability of predicting the performance from the model size.
 They also investigated the factors that affect the scaling properties of language models, including training compute, dataset size and quality, learning rate schedule, etc.
-Those experimental clues have effectively guided the later language model development, which have led to the emergence of several parameter-efficient LLMs [18, 67, 75, 1].
-However, scaling generative text-to-image models are relatively unexplored, and existing efforts have only investigated the scaling properties on small datasets or small models, like scaling UNet [46] to 270 million parameters and DiT [47] on ImageNet (14 million), or less-efficient autoregressive models [6].
+Those experimental clues have effectively guided the later language model development, which have led to the emergence of several parameter-efficient LLMs (Hoffmann et al., 2022; Touvron et al., 2023; Zhou et al., 2023; Alabdulmohsin et al., 2024).
+However, scaling generative text-to-image models are relatively unexplored, and existing efforts have only investigated the scaling properties on small datasets or small models, like scaling UNet (Nichol & Dhariwal, 2021) to 270 million parameters and DiT (Peebles & Xie, 2023) on ImageNet (14 million), or less-efficient autoregressive models (Chen et al., 2020).
 Different from these attempts, our work investigates the scaling properties by scaling down the efficient and capable diffusion models, i.e.
-LDMs [55], on internal data sources that have about 600 million aesthetics-filtered text-to-image pairs for featuring the sampling efficiency of scaled LDMs.
-We also scale LDMs on various scenarios such as finetuning LDMs on downstream tasks [69, 56] and distilling LDMs [41] for faster sampling to demonstrate the generalizability of the scaled sampling-efficiency.
+LDMs (Rombach et al., 2022), on internal data sources that have about 600 million aesthetics-filtered text-to-image pairs for featuring the sampling efficiency of scaled LDMs.
+We also scale LDMs on various scenarios such as finetuning LDMs on downstream tasks (Wang et al., 2021; Ruiz et al., 2023) and distilling LDMs (Mei et al., 2024a) for faster sampling to demonstrate the generalizability of the scaled sampling-efficiency.
 Efficient diffusion models.
 Nichol et al.
-[46] show that the generative performance of diffusion models improves as the model size increases.
-Based on this preliminary observation, the model size of widely used LDMs, e.g., Stable Diffusion [55], has been empirically increased to billions of parameters [51, 48].
+(Nichol & Dhariwal, 2021) show that the generative performance of diffusion models improves as the model size increases.
+Based on this preliminary observation, the model size of widely used LDMs, e.g., Stable Diffusion (Rombach et al., 2022), has been empirically increased to billions of parameters (Ramesh et al., 2022; Podell et al., 2023).
 However, such a large model makes it impossible to fit into the common inference budget of practical scenarios.
-Recent work on improving the sampling efficiency focus on improving network architectures [27, 74, 47, 24, 25, 8] or the sampling procedures [64, 11, 22, 34, 33, 71].
+Recent work on improving the sampling efficiency focus on improving network architectures (Li et al., 2023; Zhao et al., 2023; Peebles & Xie, 2023; Kim et al., 2023b; a; Choi et al., 2023; Mei et al., 2024b) or the sampling procedures (Song et al., 2021a; Dockhorn et al., 2022; Karras et al., 2022; Lu et al., 2022a; Liu et al., 2023c; Xu et al., 2023; Mei et al., 2025).
 We explore sampling efficiency by training smaller, more compact LDMs.
@@
 Figure 1: Text-to-image results from our scaled LDMs (39M - 2B), highlighting the improvement in visual quality with increased model size (note: 39M model is the exception).
-All images generated using 50-step DDIM sampling.
-We use representative prompts from PartiPrompts [72], including “a professional photo of a sunset behind the grand canyon.”, “Dogs sitting around a poker table with beer bottles and chips.
+All images generated using 50-step DDIM sampling and CFG rate of 7.5.
+We use representative prompts from PartiPrompts Yu et al.
+(2022), including “a professional photo of a sunset behind the grand canyon.”, “Dogs sitting around a poker table with beer bottles and chips.
 Their hands are holding cards.”, ‘Portrait of anime girl in mechanic armor in night Tokyo.”, “a teddy bear on a skateboard.”, “a pixel art corgi pizza.”, “Snow mountain and tree reflection in the lake.”, “a propaganda poster depicting a cat dressed as french emperor napoleon holding a piece of cheese.”, “a store front that has the word ‘LDMs’ written on it.”, and “ten red apples.”.
@@
 We also show a normalized running cost with respect to the baseline model.
-The text-to-image performance (FID and CLIP scores) for our scaled LDMs is evaluated on the COCO-2014 validation set with 30k samples, using 50-step DDIM sampling.
+The text-to-image performance (FID and CLIP scores) for all scaled LDMs is evaluated on the COCO-2014 validation set with 30k samples, using 50-step DDIM sampling and Classifier-free Guidance (CFG) with a rate of 7.5.
 It is worth noting that all the model sizes, and the training and the inference costs reported in this work only refer to the denoising UNet in the latent space, and do not include the 1.4B text encoder and the 250M latent encoder and decoder.
 Efficient non-diffusion generative models.
-Compared to diffusion models, other generative models such as, Variational Autoencoders (VAEs) [26, 54, 39, 68], Generative Adversarial Networks (GANs) [13, 40, 23, 52, 45], and Masked Models [10, 50, 15, 5, 4], are more efficient, as they rely less on an iterative refinement process.
+Compared to diffusion models, other generative models such as, Variational Autoencoders (VAEs) (Kingma & Welling, 2014; Rezende & Mohamed, 2015; Makhzani et al., 2015; Vahdat & Kautz, 2020), Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Mao et al., 2017; Karras et al., 2019; Reed et al., 2016; Miyato et al., 2018), and Masked Models (Devlin et al., 2019; Raffel et al., 2020; He et al., 2022; Chang et al., 2022; 2023), are more efficient, as they rely less on an iterative refinement process.
 Sauer et al.
-[60] recently scaled up StyleGAN [23] into 1 billion parameters and demonstrated the single-step GANs’ effectiveness in modeling text-to-image generation.
+(Sauer et al., 2023a) recently scaled up StyleGAN (Karras et al., 2019) into 1 billion parameters and demonstrated the single-step GANs’ effectiveness in modeling text-to-image generation.
 Chang et al.
-[4] scaled up masked transformer models for text-to-image generation.
+(Chang et al., 2023) scaled up masked transformer models for text-to-image generation.
 These non-diffusion generative models can generate high-quality images with less inference cost, which require fewer sampling steps than diffusion models and autoregressive models, but they need more parameters, i.e., 4 billion parameters.
+Figure 2: Our scaled latent diffusion models vary in the number of filters within the denoising U-Net.
+Other modules remain consistent.
+Smooth channel scaling (64 to 768) within residual blocks yields models ranging from 39M to 5B parameters.
+For downstream tasks requiring image input, we use an encoder to generate a latent code; this code is then concatenated with the noise vector in the denoising U-Net.
+Figure 3: In text-to-image generation using 50-step DDIM sampling and CFG rate of 7.5, we observe consistent trends across various model sizes in how quality metrics (FID and CLIP scores) relate to training compute (i.e., the total GFLOPS spend on training).
+Under moderate training resources, training compute is the most relevant factor dominating quality.
 3 Scaling LDMs
-We developed a family of powerful Latent Diffusion Models (LDMs) built upon the widely-used 866M Stable Diffusion v1.5 standard [55] [^1] We adopted SD v1.5 since it is among the most popular diffusion models https://huggingface.co/models?sort=likes..
+We developed a family of powerful Latent Diffusion Models (LDMs) built upon the widely-used 866M Stable Diffusion v1.5 standard (Rombach et al., 2022) [^1] We adopted SD v1.5 since it is among the most popular diffusion models https://huggingface.co/models?sort=likes..
 The denoising UNet of our models offers a flexible range of sizes, with parameters spanning from 39M to 5B.
@@
 2 shows the architectural differences during scaling.
-Models were trained using internal data sources with about 600 million aesthetically-filtered text-to-image pairs.
+Models were trained using the web-scale aesthetically filtered text-to-image dataset, i.e., WebLI (Chen et al., 2022).
 All the models are trained for 500K steps, batch size 2048, and learning rate 1e-4.
@@
 The visual quality of the results exhibits a clear improvement as model size increases.
-In order to evaluate the performance of the scaled models, we test the text-to-image performance of scaled models on the validation set of COCO 2014 [30] with 30k samples.
-For downstream performance, specifically real-world super-resolution, we test the performance of scaled models on the validation of DIV2K with 3k randomly cropped patches, which are degraded with the RealESRGAN degradation [69].
-Figure 2: Our scaled latent diffusion models vary in the number of filters within the denoising U-Net.
-Other modules remain consistent.
-Smooth channel scaling (64 to 768) within residual blocks yields models ranging from 39M to 5B parameters.
-For downstream tasks requiring image input, we use an encoder to generate a latent code; this code is then concatenated with the noise vector in the denoising U-Net.
-Refer to captionRefer to caption
-Figure 3: In text-to-image generation, we observe consistent trends across various model sizes in how quality metrics (FID and CLIP scores) relate to training compute (i.e., the product of training steps and model GFLOPS).
-Under moderate training resources ($<1G$) training compute is the most relevant factor dominating quality.
-When training compute is unlimited (e.g., $\geq 1G$), larger models ($>700M$) lead to superior performance.
-Note that the 5B model’s size exceeds single TPU capacity, forcing memory-efficient training.
-This regime has different learning dynamics than smaller models (e.g., 2B), impacting accuracy at comparable training compute.
+In order to evaluate the performance of the scaled models, we test the text-to-image performance of scaled models on the validation set of COCO 2014 (Lin et al., 2014) with 30k samples.
+For downstream performance, specifically real-world super-resolution, we test the performance of scaled models on the validation of DIV2K with 3k randomly cropped patches, which are degraded with the RealESRGAN degradation (Wang et al., 2021).
 3.1 Training compute scales text-to-image performance
-Refer to captionRefer to caption
-Figure 4: In 4×4$\times$4 × real image super-resolution, FID and LPIPS scores reveal an interesting divergence.
-Model size drives FID improvement, while training compute most impacts LPIPS.
-Despite this, visual assessment (Fig.
-5) confirms the importance of model size for superior detail recovery (similarly as observed in the text-to-image pretraining).
 We find that our scaled LDMs, across various model sizes, exhibit similar trends in generative performance relative to training compute cost, especially after training stabilizes, which typically occurs after 200K iterations.
@@
 3), the generative performance of T2I models scales well with additional compute resources.
+Figure 4: In 4×4$\times$4 × real image super-resolution using 50-step DDIM sampling, FID and LPIPS scores reveal an interesting divergence.
+Model size drives FID score improvement, while training compute most impacts LPIPS score.
+Despite this, visual assessment (Fig.
+5) confirms the importance of model size for superior detail recovery (similarly as observed in the text-to-image pretraining).
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 LR 558M 704M 866M 2B HR
-Figure 5: In 4$\times$ super-resolution, visual quality directly improves with increased model size.
+Figure 5: In 4$\times$ super-resolution using 50-step DDIM sampling, visual quality directly improves with increased model size.
 As these scaled models vary in pretraining performance, the results clearly demonstrate that pretraining boosts super-resolution capabilities in both quantitative (Fig 4) and qualitative ways.
@@
 Inputs 430M 558M 866M 2B
-Figure 6: Visualization of the Dreambooth results shows two distinct tiers based on model size.
+Figure 6: Visualization of the Dreambooth results (using 50-step DDIM sampling and CFG rate of 7.5) shows two distinct tiers based on model size.
 Smaller models (83M-223M) perform similarly, as do larger ones (318M-2B), with a clear quality advantage for the larger group.
@@
 3.2 Pretraining scales downstream performance
-Using scaled models based on their pretraining on text-to-image data, we finetune these models on the downstream tasks of real-world super-resolution [58, 57] and DreamBooth [56].
+Using scaled models based on their pretraining on text-to-image data, we finetune these models on the downstream tasks of real-world super-resolution (Saharia et al., 2022; Sahak et al., 2023) and DreamBooth (Ruiz et al., 2023).
 The performance of these pretrained models is shown in Table.
@@
 (b) 50-step sampling results of the 866M model
-Figure 7: Visualization of text-to-image results with different CFG rates (from left to right in each row: $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$).
+Figure 7: Visualization of text-to-image results with 50-step DDIM sampling and different CFG rates (from left to right in each row: $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$).
 The prompt used is “A raccoon wearing formal clothes, wearing a top hat and holding a cane.
@@
 Text-to-image generative models require nuanced evaluation beyond single metrics.
-Sampling parameters are vital for customization, with the Classifier-Free Guidance (CFG) rate [17] directly influencing the balance between visual fidelity and semantic alignment with text prompt.
+Sampling parameters are vital for customization, with the Classifier-Free Guidance (CFG) rate (Ho & Salimans, 2022) directly influencing the balance between visual fidelity and semantic alignment with text prompt.
 Rombach et al.
-[55] experimentally demonstrate that different CFG rates result in different CLIP and FID scores.
-Refer to captionRefer to caption
+(Rombach et al., 2022) experimentally demonstrate that different CFG rates result in different CLIP and FID scores.
 Figure 9: Comparison of text-to-image performance of models with varying sizes.
 The left figure shows the relationship between sampling cost (normalized cost $\times$ sampling steps) and sampling steps for different model sizes.
-The right figure plots the text-to-image FID score as a function of the sampling cost for the same models.
+The right figure plots the optimal text-to-image FID score among CFG rates of $(1.5,2.0,3.0,4.0,5.0,6.0,7.0,8.0)$ as a function of the sampling cost for the same models.
 Key Observation: Smaller models achieve better FID scores than larger models for a fixed sampling cost.
 For instance, at a cost of 3, the 83M model achieves the best FID compared to the larger models.
-This suggests that smaller models might be more efficient in achieving good visual quality with lower costs.
-We highlight the best performing model at different sampling costs [3–48].
+This suggests that smaller models can be more efficient in achieving good results with lower costs.
 Refer to caption 83M
@@
 Smaller models can produce comparable or even better visual results than larger models under similar sampling cost.
-Refer to captionRefer to caption
 Figure 11: Left: Text-to-image performance FID as a function of the sampling cost (normalized cost $\times$ sampling steps) for the DDPM sampler (solid curves) and the DDIM sampler (dashed curves).
@@
 To assess the generalizability of observed scaling trends in sampling efficiency, we compared scaled LDM performance using different diffusion samplers.
-In addition to the default DDIM sampler, we employed two representative alternatives: the stochastic DDPM sampler [16] and the high-order DPM-Solver++ [35].
+In addition to the default DDIM sampler, we employed two representative alternatives: the stochastic DDPM sampler (Ho et al., 2020) and the high-order DPM-Solver++ (Lu et al., 2022b).
 Experiments illustrated in Fig.
@@
 3.5 Scaling downstream sampling-efficiency
-Refer to captionRefer to caption
 Figure 12: Super-resolution performance vs.
@@
 Performance differences between models become less pronounced as sampling steps increase.
-Refer to captionRefer to caption
-Figure 13: Distillation improves text-to-image performance and scalability.
-Left: Distilled Latent Diffusion Models (LDMs) consistently exhibit lower (better) FID scores compared to their undistilled counterparts across varying model sizes.
-The consistent acceleration factor (approx.
-5×5$\times$5 ×) indicates that the benefits of distillation scale well with model size.
-Right: Distilled models using only 4 sampling steps achieve FID scores comparable to undistilled models using significantly more steps.
-Interestingly, at a sampling cost of 7, the distilled 866M model performs similarly to the smaller, undistilled 83M model, suggesting improved efficiency.
-See the supplement for visual examples demonstrating the quality improvements achieved through distillation.
 Here, we investigate the scaling sampling-efficiency of LDMs on downstream tasks, specifically focusing on the super-resolution task.
@@
 This observation suggests the consistent sampling efficiency of scaled models on fewer sampling steps from text-to-image generation to super-resolution tasks.
+Figure 13: Distillation improves text-to-image performance and scalability.
+Left: Distilled Latent Diffusion Models (LDMs) consistently exhibit lower (better) FID scores compared to their undistilled counterparts across varying model sizes.
+The consistent acceleration factor (approx.
+5×5$\times$5 ×) indicates that the benefits of distillation scale well with model size.
+Right: Distilled models using only 4 sampling steps achieve FID scores comparable to undistilled models using significantly more steps.
+Interestingly, at a sampling cost of 7, the distilled 866M model performs similarly to the smaller, undistilled 83M model, suggesting improved efficiency.
 3.6 Scaling sampling-efficiency in distilled LDMs.
@@
 A notable caveat, however, is that smaller models typically imply reduced modeling capability.
-This poses a challenge for recent diffusion distillation methods [36, 59, 65, 61, 14, 41, 37, 29] that heavily depend on modeling capability.
+This poses a challenge for recent diffusion distillation methods (Luhman & Luhman, 2021; Salimans & Ho, 2022; Song et al., 2023; Sauer et al., 2023b; Gu et al., 2023; Mei et al., 2024a; Luo et al., 2023; Lin et al., 2024) that heavily depend on modeling capability.
 One might expect a contradictory conclusion and believe the distilled large models sample faster than distilled small models.
-In order to demonstrate the sampling efficiency of scaled models after distillation, we distill our previously scaled models with conditional consistency distillation [65, 41] on text-to-image data and compare those distilled models on their optimal performance.
-Please see our supplement for more distillation details.
+In order to demonstrate the sampling efficiency of scaled models after distillation, we distill our previously scaled models with conditional consistency distillation (Song et al., 2023; Mei et al., 2024a) on text-to-image data and compare those distilled models on their optimal performance.
 To elaborate, we test all distilled models with the same 4-step sampling, which is shown to be able to achieve the best sampling performance; we then compare each distilled model with the undistilled one on the normalized sampling cost.
@@
 However, at the specific sampling cost, i.e., sampling cost $\approx\texttt{8}$, the smaller undistilled 83M model still achieves similar performance to the larger distilled 866M model.
-The observation further supports our proposed scaling sampling-efficiency of LDMs, which still holds under the circumstance of diffusion distillation.
+The observation further supports our proposed scaling sampling-efficiency after diffusion distillation.
 4 Conclusion
 In this paper, we investigated scaling properties of Latent Diffusion Models (LDMs), specifically through scaling model size from 39 million to 5 billion parameters.
-We trained these scaled models from scratch on a large-scale text-to-image dataset and then finetuned the pretrained models for downstream tasks.
+We trained these scaled models from scratch on a web-scale text-to-image dataset and then finetuned the pretrained models for downstream tasks.
 Our findings unveil that, under identical sampling costs, smaller models frequently outperform larger models, suggesting a promising direction for accelerating LDMs in terms of model size.
@@
 We believe this analysis of scaling sampling efficiency would be instrumental in guiding future developments of LDMs, specifically for balancing model size against performance and efficiency in a broad spectrum of practical applications.
-Limitations.
+Limitations and future work.
 This work utilizes visual quality inspection alongside established metrics like FID and CLIP scores.
 We opted to avoid human evaluations due to the immense number of different combinations needed for the more than 1000 variants considered in this study.
-However, it is important to acknowledge the potential discrepancy between visual quality and quantitative metrics, which is actively discussed in recent works [73, 19, 7].
-Furthermore, claims regarding the scalability of latent models are made specifically for the particular model family studied in this work.
-Extending this analysis to other model families, particularly those incorporating transformer-based backbones [47, 43, 38], would be a valuable direction for future research.
+However, it is important to acknowledge the potential discrepancy between visual quality and quantitative metrics, which is actively discussed in recent works (Zhang et al., 2021; Jayasumana et al., 2024; Cho et al., 2023).
+Claims regarding the scalability of latent diffusion models are made specifically for the particular model family studied in this work (Rombach et al., 2022).
+Extending this analysis to other model families, particularly those incorporating transformer-based backbones such as DiT (Peebles & Xie, 2023; Mei et al., 2023), SiT (Ma et al., 2024), MM-DiT (Esser et al., 2024), and DiS (Fei et al., 2024), and cascaded diffusion models such as Imagen3 (Baldridge et al., 2024) and Stable Cascade (Pernias et al., 2023), would be a valuable direction for future research.
 5 Acknowledgments
-We are grateful to Keren Ye, Jason Baldridge for their valuable feedback.
+Vishal M.
+Patel was supported by NSF CAREER award 2045489.
+We are grateful to Keren Ye, Jason Baldridge, Kelvin Chan for their valuable feedback.
 We also extend our gratitude to Shlomi Fruchter, Kevin Murphy, Mohammad Babaeizadeh, and Han Zhang for their instrumental contributions in facilitating the initial implementation of the latent diffusion models.
-Appendix 0.A Scaling the text-to-image performance
+Appendix A Scaling the text-to-image performance
 In order to provide detailed visual comparisons for Fig.
@@
 Figure 16: Text-to-image results from our scaled LDMs (83M - 5B), highlighting the improvement in visual quality with increased model size.
-Appendix 0.B Scaling downstream performance
+Appendix B Scaling downstream performance
 To provide more metrics for the super-resolution experiments in Fig.
@@
 5 in the main manuscript.
-Refer to captionRefer to caption
 Figure 17: For super-resolution, we show the trends between the generative metric IS and the training compute still depend on the pretraining, which is similar to the trends between the generative metric FID and the training compute.
@@
 As these scaled models vary in pretraining performance, the results clearly demonstrate that pretraining boosts super-resolution capabilities.
-Appendix 0.C Scaling sampling-efficiency in distilled LDMs
-Diffusion distillation methods for accelerating sampling are generally derived from Progressive Distillation (PD) [59] and Consistency Models (CM) [65].
-In the main paper, we have shown that CoDi [41] based on CM is scalable to different model sizes.
-Here we show other investigated methods, i.e., guided distillation [44], has inconsistent acceleration effects across different model sizes.
+Appendix C Scaling sampling-efficiency in distilled LDMs
+Diffusion distillation methods for accelerating sampling are generally derived from Progressive Distillation (PD) Salimans & Ho (2022) and Consistency Models (CM) Song et al.
+(2023).
+In the main paper, we have shown that CoDi Mei et al.
+(2024a) based on CM is scalable to different model sizes.
+Here we show other investigated methods, i.e., guided distillation Meng et al.
+(2023), has inconsistent acceleration effects across different model sizes.
 Fig.
@@
 Fig. 20 shows the visual results of the CoDi distilled models and the undistilled models under the same sampling cost to demonstrate the sampling-efficiency.
-Refer to captionRefer to caption
 Figure 19: Left: Guided distillation on the 83M model for text-to-image generation. Right: Guided distillation on the 224M model for text-to-image generation.
@@
 Figure 20: We visualize text-to-image generation results of the tested LDMs under approximately the same inference cost.
-Appendix 0.D Scaling the sampling-efficiency
+Appendix D Scaling the sampling-efficiency
 To provide more visual comparisons additional to Fig.
@@
 We observe that smaller models can produce comparable or even better visual results than larger models under similar sampling cost (model GFLOPs $\times$ sampling steps).
+Appendix E Scaling interpretability of text prompt interpolatation
+Text prompt interpolation is widely recognized as a way to evaluate the interpretability of text-to-image models in recent works (Li et al., 2024; Park et al., 2023).
+In Figure 24, we show the text-prompt interopolation results of models in different sizes and visualize their sampling results.
+Specifically, we use two distinct prompts $A$ and $B$ and interpolate their CLIP embeddings as $\alpha A+(1-\alpha)B,\alpha\in[0,1]$, to generate intermediate text-to-image results.
+A clear pattern emerges: larger models leads to more semantically coherent and visually plausible interpolations compared to their smaller counterparts.
+The figure demonstrates the 2B model’s superior ability to accurately interpret interpolated prompts, as evidenced by its generation of a tablet computer with a touch pen.
+Refer to caption 145M $\alpha=1.0$
+Refer to caption 145M $\alpha=0.8$
+Refer to caption 145M $\alpha=0.5$
+Refer to caption 145M $\alpha=0.2$
+Refer to caption 145M $\alpha=0.0$
+Refer to caption 318M $\alpha=1.0$
+Refer to caption 318M $\alpha=0.8$
+Refer to caption 318M $\alpha=0.5$
+Refer to caption 318M $\alpha=0.2$
+Refer to caption 318M $\alpha=0.0$
+Refer to caption 430M $\alpha=1.0$
+Refer to caption 430M $\alpha=0.8$
+Refer to caption 430M $\alpha=0.5$
+Refer to caption 430M $\alpha=0.2$
+Refer to caption 430M $\alpha=0.0$
+Refer to caption 866M $\alpha=1.0$
+Refer to caption 866M $\alpha=0.8$
+Refer to caption 866M $\alpha=0.5$
+Refer to caption 866M $\alpha=0.2$
+Refer to caption 866M $\alpha=0.0$
+Refer to caption 2B $\alpha=1.0$
+Refer to caption 2B $\alpha=0.8$
+Refer to caption 2B $\alpha=0.5$
+Refer to caption 2B $\alpha=0.2$
+Refer to caption 2B $\alpha=0.0$
+(a) Prompt $A$: “a old computer”. Prompt $B$: “a fancy mobile phone”.
+Figure 24: We visualize the text-prompt interpolation results of scaled models in different sizes.
+Each row shows the results of the same model with different interpolation fraction $\alpha A+(1-\alpha)B$.
+All results are sampled with the same 20-step DDIM sampler and CFG of 7.5.